# Algorithms for Optimizing Acyclic Queries

### v2

@@ 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
Before presenting the algorithm, we first formalize the relationship between $H$ and $H^{*}$ using the concept of hypergraph homomorphisms. Note the definition of homomorphisms is not standardized in the literature, and we choose ours carefully to simplify the proofs. In particular, our definition is equivalent to the homomorphism of incidence graphs [32].
@@
For two hypergraphs $H_{1}$, $H_{2}$, a homomorphism $\mathsf{f}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$ such that:
$\forall r_{1}\in R(H_{1}),x_{1}\in r_{1}:\mathsf{f}_{X}(x_{1})\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\rightarrow H_{2}$ to denote that there is a homomorphism from $H_{1}$ to $H_{2}$.
When there is no ambiguity from the context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$. Intuitively, homomorphisms preserve relations. A strong homomorphism also reflects relations:
@@
A homomorphism $\mathsf{f}:H_{1}\rightarrow H_{2}$ is called strong if it also satisfies:
$\forall r_{1}\in R(H_{1}),x_{1}^{\prime}\not\in r_{1}:\mathsf{f}_{X}(x_{1}^{\prime})\not\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\twoheadrightarrow H_{2}$ to denote that there is a strong homomorphism from $H_{1}$ to $H_{2}$.
In this paper we only consider hypergraphs over the same set of edges, [^3] The hypergraphs share the same hyperedge set $R$ but differ in the hyperedge-to-vertices functions $\chi$, i.e., an edge may have different sets of vertices in different hypergraphs. and from now on we assume $\mathsf{f}_{R}$ is the identity function.
We now show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
@@
We construct $H^{*}$ such that $H^{*}\twoheadrightarrow H$ and $H^{*}$ admits a monotonic weight join tree $\hat{T}$, by first using the MCS algorithm to construct a join tree $T$ of $H$ called the MCS tree. Then, we perturb the weights of $T$ so that they increase monotonically from the root to leaves. Finally, we construct $H^{*}$ such that the perturbed tree is a join tree of $H^{*}$.
The next result shows every MCS tree is already “somewhat monotonic”, in that every edge must contain some variable not in its parent. With abuse of notation we write $e_{1}\subseteq e_{2}$ for $\chi(e_{1})\subseteq\chi(e_{2})$ and use the common set operations directly on edges to reduce clutter.
Lemma 16.
Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
Furthermore, if two edges share any variable not in their parents, they must be siblings.
Lemma 17.
Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
We now show how to derive from an $\alpha$-acyclic hypergraph $H$ an equivalent hypergraph $H^{\prime}$ that preserves $\mathcal{T}(H)$ and admits a monotonic weight join tree $\hat{T}$.
@@
We are now ready to construct $H^{*}$ by vertex duplication.
Input: MCS tree $T_{r}(H)$
@@
2for $r^{\prime}\in c(r)$ do
3 $e\leftarrow(r,r^{\prime});\quad q$.enqueue($e$)
4
5while $q\neq\emptyset$ do
@@ 9    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
10     make $\Delta$ duplicates of $x$
11
@@
(a) $H^{*}_{6}$
(b) $\hat{T}_{p}$
Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$
Theorem 20.
Given an MCS tree $T$ of an $\alpha$-acyclic hypergraph $H$, there exists a hypergraph $H^{*}$ that
  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$ whose edges are in a one-to-one correspondence to those of $T$
  • satisfies $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
@@
To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication. Section˜4.2 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$. Algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
@@
Algorithm˜2 starts with enqueuing each edge incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. Section˜4.2 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$. By Section˜4.2, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold. The tree has monotonic weights when Algorithm˜2 terminates. ∎
Given a rooted join tree $T$ of a hypergraph $H$, a non-tree edge is $e=(r_{i},r_{j})\in E(L(H))\setminus E(T)$. There is a path in $T$ between $r_{i}$ and $r_{j}$ via their lowest common ancestor $\mathsf{LCA}(r_{i},r_{j})$. We define the LCA edges $\lambda(e)=\lambda(r_{i},r_{j})$ as a set of at most two tree edges on the path and incident to $\mathsf{LCA}(r_{i},r_{j})$. If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, Section˜4.1 and Section˜4.2 guarantee that:
  • no tree edge in $\hat{T}$ needs to be slid;
  • each non-tree edge $e$ can be removed if it is lighter than all $e^{\prime}\in\lambda(e)$;
  • otherwise, the non-tree edge $e$ can be slid directly to be incident to each $e^{\prime}\in\lambda(e)$.
Algorithm˜3 constructs $G^{\equiv}(H^{*})$ from the line graph $L(H)$ weighted by $\omega$ and an MCS tree $T_{r}$. Full details on the constant-time operations on graphs and data structures to support operations performed in Algorithm˜3 can be found in Appendix˜B. As a part of pre-processing, we first conduct a breadth-first search on $\hat{T}_{r}$ to obtain the depth table $0pt$ of each tree node in $\mathcal{O}(|R(\hat{T}_{r})|)$. We will also use the algorithms proposed by Michael Bender [3, 4] to build the data structures in $\mathcal{O}(|R(\hat{T}_{r})|)$ to facilitate the constant-time query of lowest common ancestors (LCAs) and level ancestors (LAs). The constant-time query of depth, LCA and LA allows for finding the $\lambda(e)$ of any non-tree edge $e$ in constant time as shown from Line 3 to Line 3 in Algorithm˜3.
Input: MCS tree $T_{r}(H)$ of line graph $L(H)$ with weight function $\omega$
1 $G^{\equiv}\leftarrow L(H)$
2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
 // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
3 $w^{*}=\omega(e^{*})$
4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
7 if $w^{*}<w_{1}\wedge w^{*}<w_{2}$ then
8      delete $e^{*}$ from $R(G^{\equiv})$
9
10 if $e_{1}=e_{2}\wedge w^{*}=w_{1}$ then
11      slide $r_{j}$ to $r_{1}$
12
13 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}=w_{2}$ then
14      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
15
16 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}<w_{2}$ then
17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
18
19 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{2}<w_{1}$ then
20      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
21
22return $G^{\equiv}$
Algorithm 3 buildEGbuildEG
(a) Monotonic Weight Join Tree $\hat{T}_{P}(H^{*}_{6})$
(b) 3 Non-tree Edges Processed
(c) Equivalent Graph $G^{\equiv}(H^{*}_{6})$
Figure 4: Algorithm˜3 on $H^{*}_{6}$ of Figure˜4(a)
Example 21.
We consider $\hat{T}_{P}$ in Figure˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in Figure˜4(a). Three of $\hat{T}_{P}$’s non-tree edges are shown in Figure˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed. We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time. Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed. $(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$. We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$. Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$. The intermediate graph with those three non-tree edges processed is shown in Figure˜5(b). The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in Figure˜5(c).
The next result follows immediately from the construction of the equivalent hypergraph.
@@ 6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
Section˜4.2 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges. For example, running Algorithm˜1 on the MCS tree $T(H_{6})$ in Figure˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in Figure˜5(c).
Theorem 23.
Let $H$ be an $\alpha$-acyclic hypergraph with weight function $\omega$, line graph $L$ and MCS tree $T$. Let $H^{*}$ be the equivalent hypergraph of $H$ with weight function $\omega^{*}$. Applying Algorithm˜3 to inputs $L$, $T$ and $\omega$ produces the same equivalent graph as applying Algorithm˜3 to inputs $L$, $T$ and $\omega^{*}$.
@@
One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in Algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by Section˜4.2. ∎
Theorem 24.
Given line graph $L$ of $\alpha$-acyclic hypergraph $H$ with weight function $\omega$ and an MCS tree $T$ of $H$, Algorithm˜3 returns equivalent graph $G^{\equiv}$ of equivalent hypergraph $H^{*}$ in $\mathcal{O}(m+n)$ where $n$ is the number of vertices and $m$ is the number of edges in $L$.
@@
The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [3, 4]. Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to either delete it or slide it directly to its LCA edges in the MCS tree. For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation. ∎
Once we have the equivalent graph $G^{\equiv}$, the existing algorithm allows for enumerating all its spanning trees with an amortized constant delay. The total time complexity of enumerating $|\mathcal{T}(H)|$ given the line graph $L(H)$ as input is $\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.

### v3

@@ Appendix C Missing Proofs in section˜4
We formalize the relationship between an $\alpha$-acyclic hypergraph $H$ and its equivalent hypergraph $H^{*}$ using the concept of hypergraph homomorphisms. Note the definition of homomorphisms is not standardized in the literature, [^4] A textbook on hypergraphs [5] contains two different definitions of hypergraph homomorphism; one is the same as ours, while the other is also used by Scheidt and Schweikardt [34]. and we choose ours carefully to simplify the proofs.
@@
Let $H_{1}$ and $H_{2}$ be hypergraphs. A homomorphism from $H_{1}$ to $H_{2}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$, such that for every hyperedge $r_{1}\in R(H_{1})$ and every vertex $x_{1}\in r_{1}$, the image $\mathsf{f}_{X}(x_{1})$ belongs to the hyperedge $\mathsf{f}_{R}(r_{1})$ in $H_{2}$.
We write $H_{1}\rightarrow H_{2}$ to denote the existence of such a homomorphism.
When there is no ambiguity from context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$. Intuitively, homomorphisms preserve relations. A strong homomorphism also reflects relations:
@@
Let $\mathsf{f}:H_{1}\rightarrow H_{2}$ be a hypergraph homomorphism. We call $\mathsf{f}$ strong if it also reflects non-incidence, in the following sense: for every hyperedge $r_{1}\in R(H_{1})$ and every vertex $x_{1}\in X(H_{1})$ that does not belong to $r_{1}$, the vertex $\mathsf{f}_{X}(x_{1})$ does not belong to the hyperedge $\mathsf{f}_{R}(r_{1})$ in $H_{2}$.
We write $H_{1}\twoheadrightarrow H_{2}$ to denote the existence of a strong homomorphism from $H_{1}$ to $H_{2}$.
Equivalently, under a strong homomorphism, a vertex belongs to a hyperedge in $H_{1}$ if and only if its image belongs to the image of that hyperedge in $H_{2}$.
We only consider hypergraphs over the same set of edges, [^5] The hypergraphs share the same hyperedge set $R$ but differ in incidence function $\chi$, i.e., the same edge may be incident to different vertices in different hypergraphs. and from now on we assume $\mathsf{f}_{R}$ is the identity function.
We show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
@@
Proof.
Let $T$ be a join tree of $H$, and $x^{\prime}$ be any vertex in $H^{\prime}$. Consider the neighborhood $H^{\prime}|_{x^{\prime}}$ of $x^{\prime}$ in $H^{\prime}$ comprising the set of hyperedges $\left\{r^{\prime}\in R(H^{\prime})\mid x^{\prime}\in r^{\prime}\right\}$. Since $H^{\prime}\twoheadrightarrow H$, we have that $x^{\prime}\in r^{\prime}$ if and only if $\mathsf{f}(x^{\prime})\in\mathsf{f}(r^{\prime})=r^{\prime}$. The neighborhood $H|_{\mathsf{f}(x^{\prime})}$ of $\mathsf{f}(x^{\prime})$ in $H$ comprises of the set of hyperedges $\left\{r\in R(H)\mid\mathsf{f}(x^{\prime})\in r\right\}$. Therefore, $r^{\prime}\in R(H^{\prime}|_{x^{\prime}})\iff\mathsf{f}(r^{\prime})=r^{\prime}\in R(H|_{\mathsf{f}(x^{\prime})})$. Because $T$ is a join tree of $H$, $H|_{\mathsf{f}(x^{\prime})}$ is connected in $T$, therefore $H^{\prime}|_{x^{\prime}}$ is also connected in $T$ and $T$ is also a join tree of $H^{\prime}$. ∎
Our goal will be to modify an input hypergraph $H$ such that its line graph admits a monotonic weight join tree. The lemma above ensures that if this is done in such a way that a strong homomorphism between $H$ and its modification is preserved then then the set of join trees of $H$ is preserved.
@@
Proof.
2. It follows from section˜C.1 and the facts above.
Input: $\alpha$-acyclic hypergraph $H$ and rooted MCS tree $T_{r}$
Output: $H^{*}$ in Theorem˜21
@@
[…]
4for $r^{\prime}\in c(r)$ in $T_{r}$ do
5 $e\leftarrow\{r,r^{\prime}\};\quad q$.enqueue($e$)
6
7while $q\neq\emptyset$ do
@@ 11    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
12     make $\Delta$ duplicates of $x$ in $X^{*}$
13
@@
Vertex duplication allows us to modify a hypergraph while preserving its set of join trees. In particular, duplicating vertices can selectively increase the weights of chosen hyperedge intersections without affecting the line graph or the join tree structure (cf. section˜C.1).
We use this operation to enforce monotonicity of edge weights along a given join tree. Given an $\alpha$-acyclic hypergraph $H$ together with a rooted MCS tree $T$, we can construct an equivalent hypergraph $H^{*}$ by duplicating vertices along the edges of $T$. The construction proceeds in a breadth-first manner from the root of $T$ and duplicates vertices only when necessary to ensure that each tree edge is strictly heavier than its parent.
Obtaining an equivalent graph in this manner always succeeds.
Theorem˜21 Given an $\alpha$-acyclic hypergraph $H=(X,R,\chi)$, there exists an equivalent hypergraph $H^{*}=(X^{*},R,\chi^{*})$ over the same hyperedge set $R$ that
  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$, and
  • $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
@@
To prove existence, we construct hypergraph $H^{*}$ by vertex duplication (algorithm˜6). section˜C.1 will guarantee that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. Therefore, we only need to show the existence of monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$. algorithm˜6 takes as input hypergraph $H$ along with an MCS tree $T$ of $H$. The algorithm constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
@@
algorithm˜6 starts with enqueuing each edge of $T$ incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. section˜3 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
We prove that this produces a monotonic weight join tree by induction on a breadth-first ordering (with respect to root $r$) of the edges in $T$.
Base case: The base case is an edge $e$ incident to the root. Since $e$ has no parent $\omega(e)>\omega(p(e))$ holds vacuously.
Ind. step: Let $(e_{1},\ldots,e_{m})$ be a breadth-first ordering of the edges in $T$ and assume $\omega(e_{i})>\omega(p(e_{i}))$ for all i<k1i<k-1.
Let $e_{k}=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$ and let $x\in\chi(e_{k})\setminus\chi(p(e_{k}))$. Vertex $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge $e_{i}$, $i\in[k-1]$ outside of subtree rooted at $r_{u}$ and therefore $\omega(e_{i})>\omega(p(e_{i}))$ holds under any number of duplications of $x$. For any $e_{i}$ lying within the subtree rooted at $r_{u}$, $e_{i}$ is necessarily a sibling of $e_{k}$ since edges of $T$ are breadth-first ordered. Hence $p(e_{i})$ lies outside this subtree and any number of vertex duplications of $x$ cannot decrease the weight of $e_{i}$ with respect to its parent. Therefore, $\omega(e_{i})>\omega(p(e_{i}))$ for all $i\in[k]$ after performing $\Delta$ many vertex duplications of $x$.
Consequently, tree $T$ has monotonic weights when algorithm˜6 terminates. ∎
The above theorem guarantees that for any $\alpha$-acyclic hypergraph we are able to obtain an equivalent hypergraph $H^{*}$. Next we prove Theorem˜24 showing that algorithm˜2 produces the equivalent graph of line graph $L(H^{*})$.
Theorem˜24 Given a hypergraph $H^{*}=(X^{*},R,\chi^{*})$ with a monotonic weight join tree $\hat{T}$, algorithm˜2 produces an Equivalent Graph $G^{\equiv}(L(H^{*}))$.
Proof.
[…]
Of note is the case where a non-tree edge $e$ weighs less than all edges in $\lambda(e)$. The algorithm then slides $e$ to a self loop at $l=\mathsf{LCA}(e)$. Such an edge may still be able to slide if edge $\{l,p(l)\}$ weighs more than edge $e$. In this sense graph $G$ returned by algorithm˜2 may not be maximal under partial order $\leq_{\hat{T}}$. Therefore it remains to show that these self loops do not prevent $G$ from being an equivalent graph. To see this, notice that no self loop can take part in a spanning tree. Therefore any spanning tree of graph obtained upon sliding self loops of $G$ to a fixpoint ($G^{\equiv}$) is also a spanning tree of $G$. Hence, $G$ and $G^{\equiv}$ have the same set of spanning trees and $G$ is an equivalent graph.
The next result follows from the construction of the equivalent hypergraph.
@@
As section˜4 shows, running algorithm˜2 on the MCS tree $T(H_{6})$ in fig.˜2(d) produces $G^{\equiv}(H_{6})$ identical to $G^{\equiv}(H^{*}_{6})$ in fig.˜5(c). algorithm˜2 does this in time linear in the size of the line graph. We give the formal proof below.
Theorem˜26 Given the line graph $L$ and an MCS tree $T_{r}$ of an $\alpha$-acyclic hypergraph $H$, algorithm˜2 returns an equivalent graph $G^{\equiv}(L)$ in time $\mathcal{O}(|L|)$.
@@
Let $H^{*}=(X^{*},R,\chi^{*})$ be an equivalent hypergraph derived from $H=(X,R,\chi)$ by vertex duplication. One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in algorithm˜2 hold under weight function $\omega$ of $L(H)$ if and only if they hold under weight function $\omega^{*}$ of $L(H^{*})$ by section˜C.1. Therefore algorithm˜2 returns an equivalent graph $G^{\equiv}(L)$.
The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [4, 3]. Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to slide it directly to its LCA edges in the MCS tree. For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation. The overall runtime is thus $\mathcal{O}(|L|)$. ∎
Given algorithm˜2, we now have a systematic way of enumerating join trees of an $\alpha$-acyclic hypergraph.
Starting from the line graph $L(H)$ of an $\alpha$-acyclic hypergraph $H$, we first reconstruct $H$ and apply algorithm˜1 to obtain an MCS tree $T$. Using $L(H)$ and $T$ as input, algorithm˜2 constructs an equivalent graph of $L(H)$. Finally, a standard spanning tree enumeration algorithm can be applied to this equivalent graph to enumerate all join trees of $H$. The overall running time of this process is $\mathcal{O}(|L|+|\mathcal{T}(H)|)$.
Theorem˜27 Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, the join trees of $H$ can be enumerated in time $\mathcal{O}(|L|+|\mathcal{T}(H)|)$.